Compute the total and annual returns on the following investment.

Fifteen years after purchasing shares in a mutual fund for ​$6200 the shares are sold for $11,500

1-the total return is _% (rounded to one decimal)

2-The annual return is _%
Answer #1

Hi

Here initial Value of investment P = $6200

Sold Value F = $11,500

1) Total Return = F/P-1

= 11500/6200 - 1

= 85.48%

2) Annual return = (F/P)^(1/15) -1

= (11500/6200)^(1/15) -1

= 4.20%

Thanks